How many gallons fo a 20% acid solution should be mixed with a 30% acid solution to produce 100 gallons of a 23% solution?
Found 2 solutions by ewatrrr, richwmiller:
Answer by ewatrrr(24785)

Set Up
x 30% amount
.30x + .20(100G - x) = .23(100G)
x = .03(100G)/.10 = 30G, the 30% amount. 70G of the 20% 100-30

Answer by richwmiller(17219)

a+b=100,
0.2*a+0.3*b=0.23*100
a=100-b
0.2*(100-b)+0.3*b=23
20-0.2b+0.3*b=23
0.1*b=3
b=30
a=100-b
a=70 gal at 20% b=30 gal at 30%
check
0.2*70+0.3*30=0.23*100
14+9=23
23=23
ok